OpenStudy (fantalux):

Which ordered pair in the form (x, y) is a solution of this equation? (x + 3)y = 14 A. (3, 2) B. (7, 2) C. (11, 1) D. (5, 2) Which pairs are solutions to the equation? 3xy + 9 = 33 A. (1, 11) and (2, 6) B. (8, 1) and (2, 6) C. (1, 8) and (1, 11) D. (1, 8) and (8, 1)

OpenStudy (anonymous):

In these type of question, the easiest way to get the answer is try to plug in the x and y values from the options given in the equation given, So in the first question all the choice except C are more then 14 if you plug in x and y's, for eg, if you plug in x = 3 and y = 2 , you get (3+3)2 = 14 6 x 2 = 14 12 is not equal to 14, so this eliminates this choice but if you chose C you get, (11+3)1 = 14 14 = 14 so this makes C the solution for first question and for the second question do the same thing, and the answer will be D. Hope this helps

OpenStudy (fantalux):

My Guess was that 1 is C and 2 is D......Am i correct?

OpenStudy (fantalux):

oh oops just read your whole responce thanks..

OpenStudy (anonymous):

yes you are, well if you wanna make sure if you are right or wrong just plug the x and y from the options like i said and if the left hand side of the equation is equal to the right hand side (for eg, 14 = 14) then it is correct.

OpenStudy (fantalux):

Hey could you help me with these 5 other questions someone else tried to help me but they were all wrong :(

OpenStudy (fantalux):

I just overlooked the 5 questions and there is only a couple i need help with....

OpenStudy (fantalux):

Solve the equation for y. 5x – 7y = 10